# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ II-B Operating System (OS) of LM Agents
According to [47, 48], the OS architecture of LM agents consists of three layers: application, kernel, and hardware, as shown in Fig. 5.
@@
  • The kernel layer includes the ordinary OS kernel (to process LM-related queries such as file operation and network request) and an additional LM agent kernel (to process LM-related queries such as reasoning, planning, and tool parsing), with a focus on without altering the original OS structure. Key modules in the LM agent kernel [47, 48] include the agent scheduler for task planning and prioritization, context manager for LM status management, memory manager for short-term data, storage manager for long-term data retention, tool manager for external API interactions, and access manager for privacy controls.
@@ II-C Constructing Modules of LM Agents
According to [9, 2, 10, 11], there are generally five constructing modules of LM agents: planning, action, memory, interaction, and security. As depicted in Fig. 6, these modules together enable LM agents to perceive, plan, act, learn, and interact efficiently and securely in complex and dynamic environments. Particularly, empowered by LMs, the planning module produces strategies and action plans with the help of the memory module, enabling informed decision-making. The action module executes these embodied actions, adapting actions based on real-time environmental feedback to ensure contextually appropriate responses. The memory module serves as a repository of accumulated knowledge (e.g., past experiences and external knowledge), facilitating continuous learning and improvement. The interaction module enables effective communication and collaboration with humans, other agents, and environment. The security module is integrated throughout LM agents’ operations, ensuring active protection against threats and maintaining integrity and confidentiality of data and processes.
@@ II-C1 Planning Module
The planning module serves as the core of an LM agent [8, 11]. It utilizes advanced reasoning techniques to devise effective solutions fo agents to complex problems. It includes the following working modes.
  • Feedback-free planning: The planning module enables LM agents to understand the complex problems and find reliable solutions by breaking them down into necessary steps or manageable sub-tasks [8, 12], as shown in Fig. 8. For instance, chain-of-thought (CoT) [49] is a popular sequential reasoning approach where each thought builds directly on the previous one. It represents the step-by-step logical thinking and can enhance the generation of coherent and contextually relevant responses. Tree-of-thought (ToT) [50] organizes reasoning as a tree-like structure, exploring multiple paths simultaneously. In ToT, each node represents a partial solution, allowing the model to branch and backtrack to find the optimal answer. Graph-of-thought (GoT) [51] models reasoning using an arbitrary graph structure, allowing more flexible information flow. GoT captures complex relationships between thoughts, enhancing the model’s problem-solving capabilities. AVIS [55] further refines the tree search process for visual QA tasks using a human-defined transition graph and enhances decision-making through a dynamic prompt manager.
  • Feedback-enhanced planning: To make effective long-term planning in complex tasks, it is necessary to iteratively reflect on and refine execution plans based on past actions and observations [41]. The goal is to correct past errors and improve final outcomes. For instance, as shown in Fig. 9, ReAct [52] combines reasoning and acting by prompting LLMs to generate reasoning traces (i.e., thoughts) and actions simultaneously. This dual approach allows the LLM to create, monitor, and adjust action plans, while task-specific actions enhance interaction with external sources, improving response accuracy and reliability. As shown in Fig. 10(a), Reflexion [53] converts environmental feedback into self-reflection and enhances ReAct by enabling LLM agents to learn from past errors and iteratively optimize behaviors. Reflexion features an actor that produces actions and text via models (e.g., CoT and ReAct) enhanced by memory, an evaluator that scores outputs using task-specific reward functions, and self-reflection that generates verbal feedback to improve the actor.
@@
  • Grounded planning: Executing plans in real or simulated world environments (e.g., Minecraft) requires precise, multi-step reasoning. As shown in Fig. 10(b), VOYAGER [54], the first LLM-powered agent in Minecraft, utilizes in-context lifelong learning to adapt and generalize skills to new tasks and worlds. VOYAGER include an automatic curriculum for exploration, a skill library of executable code for complex behaviors, and an iterative prompting mechanism that refines programs based on feedback. Wang et al. [57] further propose an interactive describe-explain-plan-select (DEPS) planning approach that improves LLM-generated plans by integrating execution descriptions, self-explanations, and a goal selector that ranks sub-goals to refine planning.
@@ II-C2 Memory Module
  • Short-term memory focuses on the contextual information of the current situation. It is temporary and limited, typically managed through a context window that restricts the amount of information the LM agent can learn at a time.
@@
  • Hybrid memory combines short-term and long-term memory to enhance an agent’s understanding of the current context and leverage past experiences for better long-term reasoning. Liu et al. [60] propose the RAISE architecture to enhance ReAct for conversational AI agents by integrating a dual-component memory system, where Scratchpad captures recent interactions as short-term memory; while the retrieval module acts as long-term memory to access relevant examples.
@@ II-C3 Action Module
The action module equips the LM agent with the ability to execute and adapt actions in various environments [10, 45]. This module is designed to handle embodied actions and tool-use capabilities, enabling the agent to interact with its physical surroundings adaptively and effectively. Besides, tools significantly broaden the action space of the agent.
  • Embodied actions. The action module empowers LM agents to perform contextually appropriate embodied actions and adapt to environmental changes, facilitating interaction with and adjustment to physical surroundings [61, 15]. As LLM-generated action plans are often not directly executable in interactive environments, Huang et al. [61] propose refining LLM-generated plans for embodied agents by conditioning on demonstrations and semantically translating them into admissible actions. Evaluations in the VirtualHome environment show significant improvements in executability, ranging from 18% to 79% over the baseline LLM. Besides, SayCan [15] enables embodied agents such as robots to follow high-level instructions by leveraging LLM knowledge in physically grounded tasks, where LLM (i.e., Say) suggests useful actions; while learned affordance functions (i.e., Can) assess feasibility. SayCan’s effectiveness is demonstrated through 101 zero-shot real-world robotic tasks in a kitchen setting. PaLM-E [62] is a versatile multimodal language model for embodied reasoning, visual-language, and language tasks. It integrates continuous sensor inputs, e.g., images and state estimates, into the same embedding space as language tokens, allowing for grounded inferences in real-world sequential decision-making.
  • Learning to use & make tools. By leveraging various tools (e.g., search engines and external APIs) [63], LM agents can gather valuable information to handle assigned complex tasks. For instance, AutoGPT integrates LLMs with predetermined tools such as web and file browsing. InteRecAgent [17] integrates LLMs as the brain and recommender models as tools, using querying, retrieval, and ranking tools to handle complex user inquiries. Beyond using existing tools, LM agents can also develop new tools to enhance task efficiency [10]. To optimize tool selection with a large toolset, ReInvoke [14] introduces an unsupervised tool retrieval method featuring a query generator to enrich tool documents in offline indexing and an intent extractor to identify tool-related intents from user queries in online inference, followed by a multi-view similarity ranking strategy to identify the most relevant tools.
@@ II-C4 Interaction Module
  • Agent-Environment Interactions. LM agents can engage directly with the physical or virtual environments. By facilitating engagement in physical, virtual, or mixed-reality environments [15, 2], the interaction module ensures that LM agents can operate effectively across different contexts. Lai et al. develop the AutoWebGLM agent [66], which excels in web browsing tasks through curriculum learning, self-sampling reinforcement learning, and rejection sampling fine-tuning. A Chrome extension based on AutoWebGLM validates its effective reasoning and operation capability across various websites in real-world services.
@@ II-C5 Security Module
Connections Between Modules: The above five key components of an LM agent are interconnected to create a cohesive and intelligent system. Particularly, the planning module relies on the memory module to access past experiences and external knowledge, ensuring informed decision-making. The action module executes plans generated by the planning module, adapting actions based on real-time feedback and memory. The interaction module enhances these processes by facilitating communication and collaboration, which provides additional data and context for the planning and memory modules. Besides, security considerations are seamlessly integrated into every aspect of the LM agent’s operations to ensure robust and trusted executions.
TABLE III: Comparison of Existing Typical LM Agent Prototypes
@@
Other Key Features
AutoGPT General scenarios Yes N/A Yes Task decomposition and execution
AutoGen General scenarios Yes N/A Yes Multi-agent conversation framework
BabyAGI General scenarios Yes N/A Yes Autonomous task management simulation
ChatDev General scenarios Yes N/A Yes Conversational AI for software development
MetaGPT General scenarios Yes N/A Yes Meta-learning for task adaptation
MobileAgent v2 General scenarios Yes N/A Yes Multimodal input support
AppAgent General scenarios Yes N/A Yes Performs tasks on mobile devices
NetLLM Mobile communications Yes N/A Yes Fine-tunes LLM for networking
NetGPT Mobile communications Yes N/A Yes Cloud-edge cooperative LM service
Figure 02 by FigureAI Intelligent robots Yes Yes N/A Performs dangerous jobs
Optimus by Tesla Intelligent robots Yes Yes N/A Second-generation humanoid robot
Baidu Apollo ADFM Autonomous driving Yes Yes Yes Supports L4 autonomous driving
PentestGPT Attack-defense confrontation Yes Yes N/A 87% success in vulnerability exploitation
AutoAttacker Attack-defense confrontation Yes No N/A Automatically execute network attacks
@@ II-D1 General Scenarios
Recently, various industrial projects of LM agents, such as AutoGPT, AutoGen, BabyAGI, ChatDev, and MetaGPT, demonstrate their diverse potential in assisting web, life, and business scenarios, such as planning personalized travels, automating creative content generation, and enhancing software development workflows. For instance, AutoGPT [^3] https://autogpt.net/ is an open-source autonomous agent utilizing GPT-3.5 or GPT-4 APIs to independently execute complex tasks by breaking down them into several sub-tasks and chaining LLM outputs, showcasing advanced reasoning capabilities [12]. AutoGen [^4] https://microsoft.github.io/autogen/, developed by Microsoft, offers an open-source multi-agent conversation framework, supports APIs as tools for improved LLM inference, and emphasizes the automatic generation and fine-tuning of AI models [13]. BabyAGI [^5] https://github.com/yoheinakajima/babyagi integrates task management via OpenAI platforms and vector databases, simulating a simplified AGI by autonomously creating and executing tasks based on high-level objectives. ChatDev [^6] https://github.com/OpenBMB/ChatDev focuses on enhancing conversational AI, providing sophisticated dialogue management, coding, debugging, and project management to streamline software development processes [67]. MetaGPT [^7] https://www.deepwisdom.ai/ explores the meta-learning paradigm, where the model is trained to rapidly adapt to new tasks by leveraging knowledge from related tasks, thus improving performance across diverse applications [68].
LM agents can also enhance the QoE of end users. For instance, MobileAgent v2 [18], launched by Alibaba, is a mobile device operation assistant that achieves effective navigation through multi-agent collaboration, automatically performing tasks such as map navigation, and supports multimodal input including visual perception, enhancing operational efficiency on mobile devices. AppAgent [19], developed by Tencent, performs various tasks on mobile phones through autonomous learning and imitating human click and swipe gestures, including posting on social media, helping users write and send emails, using maps, online shopping, and even complex image editing.
@@ II-D2 Mobile Communications
LM agents offer significant advantages for mobile communications by enabling low-cost and context-aware decision-making [69], personalized user experiences [70], and automatic optimization problem formulation for wireless resource allocation [71]. For instance, NetLLM [69] fine-tunes the LLM to acquire domain knowledge from multimodal data in networking scenarios (e.g., adaptive bitrate streaming, viewport prediction, and cluster job scheduling) with reduced handcraft costs. Meanwhile, NetGPT [70] uses a cloud-edge cooperative LM framework for personalized outputs and enhanced prompt responses in mobile communications via de-duplication and prompt enhancement technologies. ChatNet [72] uses four GPT-4 models to serve as analyzer (to plan network capacity and designate tools), planner (to decouple network tasks), calculator (to compute and optimize the cost), and executor (to produce customized network capacity solutions) via prompt engineering.
@@ II-D3 Intelligent Robots
LM agents play a crucial role in advancing intelligent industrial and service robots [15]. These robots can perform complex tasks such as product assembly, environmental cleaning, and customer service, by perceiving surroundings and learning necessary skills through LMs. In August 2024, FigureAI released Figure 02888https://spectrum.ieee.org/figure-new-humanoid-robot, a human-like robot powered by OpenAI LM, capable of fast common-sense visual reasoning and speech-to-speech conversation with humans to handle dangerous jobs in various environments. Besides, in July 2024, Tesla unveiled its second-generation humanoid robot named Optimus [^9] https://viso.ai/edge-ai/tesla-bot-optimus/, which demonstrates the enhanced capabilities brought by LM agents.
@@ II-D4 Autonomous Driving
LM agents are transforming autonomous driving by enhancing vehicle intelligence, improving safety, and optimizing driving experience (e.g., offering personalized in-car experience) [21]. In May 2024, Baidu’s Apollo Day saw the launch of Carrot Run’s sixth-generation unmanned vehicle, which is built upon the Apollo autonomous driving foundation model (ADFM) [^10] https://autonews.gasgoo.com/icv/70033042.html, the first LM agent supporting L4 autonomous driving. Companies such as Tesla, Waymo, and Cruise are also leveraging LM agents to refine their autonomous driving systems, aiming for safer and more efficient transportation systems.
Figure 11: General architecture of LM agent networks in bridging the human, physical, and cyber worlds. The LM agent has five constructing modules: planning, action, memory, interaction, and security modules. The engine of connected LM agents is empowered by a combination of five cutting-edge technologies: foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration. LM agents interact with humans through HMI technologies such as NLP, with the assistance of handheld devices and wearables, to understand human intentions, desires, and beliefs. LM agents can synchronize data and statuses between the physical body and the digital brain through digital twin technologies, as well as perceive and act upon the surrounding virtual/real environment. LM agents can be interconnected in the cyber space through efficient cloud-edge networking for efficient sharing of data, knowledge, and computation results to promote multi-agent collaboration.

### v3

@@ II-B Operating System (OS) of LM Agents
According to [43], the OS architecture of LM agents consists of three layers: application, kernel, and hardware, as shown in Fig. 6.
@@
  • The kernel layer includes the ordinary OS kernel (to process LM-related queries such as file operation and network request) and an additional LM agent kernel (to process LM-related queries such as reasoning, planning, and tool parsing), with a focus on without altering the original OS structure. Key modules in the LM agent kernel [43] include the agent scheduler for task planning and prioritization, context manager for LM status management, memory manager for short-term data, storage manager for long-term data retention, tool manager for external API interactions, and access manager for privacy controls.
@@ II-C Constructing Modules of LM Agents
According to [2, 7], there are generally five constructing modules of LM agents: planning, action, memory, interaction, and security. As depicted in Fig. 7, these modules together enable LM agents to perceive, plan, act, learn, and interact efficiently and securely in complex and dynamic environments. Particularly, empowered by LMs, the planning module produces strategies and action plans with the help of the memory module, enabling informed decision-making. The action module executes these embodied actions, adapting actions based on real-time environmental feedback to ensure contextually appropriate responses. The memory module serves as a repository of accumulated knowledge (e.g., past experiences and external knowledge), facilitating continuous learning and improvement. The interaction module enables effective communication and collaboration with humans, other agents, and environment. The security module is integrated throughout LM agents’ operations, ensuring active protection against threats and maintaining integrity and confidentiality of data and processes.
@@ II-C1 Planning Module
The planning module serves as the core of an LM agent [6, 7]. It utilizes advanced reasoning techniques to devise effective solutions for agents to complex problems. It includes the following working modes.
  • Feedback-free planning: The planning module enables LM agents to understand the complex problems and find reliable solutions by breaking them down into necessary steps or manageable sub-tasks [6, 8], as shown in Fig. 9. For instance, chain-of-thought (CoT) [44] is a popular sequential reasoning approach where each thought builds directly on the previous one. It represents the step-by-step logical thinking and can enhance the generation of coherent and contextually relevant responses. Tree-of-thought (ToT) [45] organizes reasoning as a tree-like structure, exploring multiple paths simultaneously. In ToT, each node represents a partial solution, allowing the model to branch and backtrack to find the optimal answer. Graph-of-thought (GoT) [46] models reasoning using an arbitrary graph structure, allowing more flexible information flow. GoT captures complex relationships between thoughts, enhancing the model’s problem-solving capabilities. AVIS [50] further refines the tree search process for visual QA tasks by leveraging a human-defined transition graph and improves decision-making through a dynamic prompt manager.
  • Feedback-enhanced planning: To make effective long-term planning in complex tasks, it is necessary to iteratively reflect on and refine execution plans based on past actions and observations [35]. The goal is to correct past errors and improve final results. For instance, as illustrated in Fig. 10, ReAct [47] integrates reasoning and action by prompting LLMs to concurrently produce reasoning traces (i.e., thoughts) and actions. This dual approach allows the LLM to create, monitor, and adjust action plans, while task-specific actions enhance interaction with external sources, improving response accuracy and reliability. As shown in Fig. 11(a), Reflexion [48] converts environmental feedback into self-reflection and enhances ReAct by enabling LLM agents to learn from past errors and iteratively optimize behaviors. Reflexion features an actor that produces actions and text via models (e.g., CoT and ReAct) enhanced by memory, an evaluator that scores outputs using task-specific reward functions, and self-reflection that generates verbal feedback to improve the actor.
@@
  • Grounded planning: Executing plans in real or simulated world environments (e.g., Minecraft) requires precise, multi-step reasoning. As shown in Fig. 11(b), VOYAGER [49], the first LLM-powered agent in Minecraft, utilizes in-context lifelong learning to adapt and generalize skills to new tasks and worlds. VOYAGER incorporates an automated curriculum for exploration, a skill library containing executable code for complex behaviors, and an iterative prompting mechanism that refines programs based on feedback. Wang et al. [52] further propose an interactive describe-explain-plan-select (DEPS) planning approach that improves LLM-generated plans by integrating execution descriptions, self-explanations, and a goal selector that ranks sub-goals to refine planning.
@@ II-C2 Memory Module
  • Short-term memory focuses on the contextual information of the current situation. It is temporary and limited, typically managed through a context window that restricts the amount of information the LM agent can learn at a time [54].
@@
  • Hybrid memory synergizes short-term and long-term memory to enhance an agent’s understanding of the current context and leverage past experiences for better long-term reasoning. Liu et al. [56] propose the RAISE architecture to enhance ReAct for conversational AI agents by integrating a dual-component memory system, where Scratchpad captures recent interactions as short-term memory; while the retrieval module acts as long-term memory to access relevant examples.
@@ II-C3 Action Module
The action module equips the LM agent with the ability to execute and adapt actions in various environments [2, 38]. This module is designed to handle embodied actions and tool-use capabilities, allowing the agent to interact with its physical surroundings adaptively and effectively. Besides, tools significantly broaden the agent’s action space.
  • Embodied actions. The action module empowers LM agents to perform contextually appropriate embodied actions and adapt to environmental changes, facilitating interaction with and adjustment to physical surroundings [57, 11]. As LLM-generated action plans are often not directly executable in interactive environments, Huang et al. [57] propose refining LLM-generated plans for embodied agents by conditioning on demonstrations and semantically translating them into admissible actions. Evaluations in the VirtualHome environment show significant improvements in executability, ranging from 18% to 79% over the baseline LLM. Besides, SayCan [11] enables embodied agents such as robots to follow high-level instructions by leveraging LLM knowledge in physically grounded tasks, where LLM (i.e., Say) suggests useful actions; while learned affordance functions (i.e., Can) assess feasibility. SayCan’s effectiveness is demonstrated through 101 zero-shot real-world robotic tasks in a kitchen setting. PaLM-E [58] is a versatile multimodal language model for embodied visual-language and language tasks. It integrates continuous sensor inputs, e.g., images and state estimates, into the same embedding space as language tokens, allowing for grounded inferences in real-world sequential decision-making.
  • Learning to use & make tools. By leveraging various tools (e.g., search engines and external APIs) [59], LM agents can gather valuable information to handle assigned complex tasks. For instance, AutoGPT integrates LLMs with predetermined tools such as web and file browsing. Beyond using existing tools, LM agents can also develop new tools to enhance task efficiency [2]. To optimize tool selection with a large toolset, ReInvoke [10] introduces an unsupervised tool retrieval method featuring a query generator to enrich tool documents in offline indexing and an intent extractor to identify tool-related intents from user queries in online inference, followed by a multi-view similarity ranking strategy to identify the most relevant tools.
Figure 12: The taxonomy of type and constructing modules of LM agents.
@@ II-C4 Interaction Module
  • Agent-Environment Interactions. LM agents can engage directly with the physical or virtual environments. By facilitating engagement in physical, virtual, or mixed-reality environments [11], the interaction module ensures that LM agents can operate effectively across different contexts. Lai et al. develop the AutoWebGLM agent [62], which excels in web browsing missions via self-sampling reinforcement learning, curriculum learning, and rejection sampling fine-tuning. A Chrome extension based on AutoWebGLM validates its effective reasoning and operation capability across various websites in real-world services.
@@ II-C5 Security Module
Connections Between Modules: The above five key components of an LM agent are interconnected to create a cohesive and intelligent system. Particularly, the planning module relies on the memory module to access past experiences and external knowledge, ensuring informed decision-making. The action module executes plans generated by the planning module, adapting actions based on real-time feedback and memory. The interaction module enhances these processes by facilitating communication and collaboration, which provides additional data and context for the planning and memory modules. Besides, security considerations are seamlessly integrated into every aspect of the LM agent’s operations to ensure robust and trusted executions. Fig. 12 shows the taxonomy of type and key modules of LM agents.
TABLE IV: Comparison of Typical LM Agent Prototypes
@@
Other Key Features Limitations
AutoGPT[8]
General
scenarios
Yes N/A Yes
Task decomposition and
execution
Ecosystem isolation,
single-device simulation
AutoGen[9]
General
scenarios
Yes N/A Yes
Multi-agent conversation
framework
Ecosystem isolation,
single-device simulation
ChatDev[63]
General
scenarios
Yes N/A Yes
Conversational AI for
software development
Ecosystem isolation,
single-device simulation
MetaGPT[64]
General
scenarios
Yes N/A Yes
Meta-learning for task
adaptation
Ecosystem isolation,
single-device simulation
MobileAgent v2[13]
General
scenarios
Yes N/A Yes
Multimodal input support
Unstable performance
for languages
NetLLM[65]
Mobile
communications
Yes N/A Yes
Fine-tunes LLM for
networking scenarios
Suboptimal performance
of multimodal fusion
NetGPT[66]
Mobile
communications
Yes N/A Yes
Cloud-edge cooperative
LM service
Missing exploration of
multimodal LMs
Figure 02[67] Intelligent robots Yes Yes N/A Performs dangerous jobs
Simulation-to-reality gap
in training
Unitree G1[68] Intelligent robots Yes Yes N/A
2nd generation humanoid
robot
Limited autonomy in
complex environments
Apollo ADFM[69]
Autonomous
driving
Yes Yes Yes
Supports L4 autonomous
driving
Challenges in handling
extreme edge cases
PentestGPT[16]
Attack-defense
confrontation
Yes Yes N/A
87% success in
vulnerability exploitation
Limited effectiveness in
high-complexity scenes
AutoAttacker[70]
Attack-defense
confrontation
Yes No N/A
Automatically execute
network attacks
Narrow scope,
LLM reliability risks
@@ II-D1 General Scenarios
Recently, various industrial projects of LM agents, such as AutoGPT, AutoGen, BabyAGI, ChatDev, and MetaGPT, demonstrate their diverse potential in assisting web, life, and business scenarios, such as planning personalized travels, automating creative content generation, and enhancing software development workflows. For instance, AutoGPT [8] is an open-source autonomous agent utilizing GPT-3.5 or GPT-4 APIs to independently execute complex tasks by breaking down them into several sub-tasks and chaining LLM outputs, showcasing advanced reasoning capabilities. AutoGen [9], developed by Microsoft, offers an open-source multi-agent conversation framework, supports APIs as tools for improved LLM inference, and emphasizes the automatic generation and fine-tuning of AI models. ChatDev [63] focuses on enhancing conversational AI, providing sophisticated dialogue management, coding, debugging, and project management to streamline software development processes. MetaGPT [64] explores the meta-learning paradigm, where the model is trained to rapidly adapt to new tasks by leveraging knowledge from related tasks, thus improving performance across diverse applications.
LM agents can also enhance the QoE of end users. For instance, MobileAgent v2 [13], launched by Alibaba, is a mobile device operation assistant that achieves effective navigation through multi-agent collaboration, automatically performing tasks such as map navigation, and supports multimodal input including visual perception, enhancing operational efficiency on mobile devices. Besides, DeepSeek-R1 [71] has been widely adopted across industries, including finance, education, healthcare, office automation, and AI assistants. Leading companies such as Baidu, AWS, Azure AI, and Qihoo 360 have integrated DeepSeek-R1 into their applications. Additionally, AI agent platforms such as Coze AI and BetterYeah AI enable seamless agent deployment and customization using DeepSeek-R1. For instance, AgenticFlow [72] automates marketing campaigns through AI-driven workflows, while Dify [73] leverages DeepSeek-R1 to build assistants, workflows, and text generators. RAGFlow [74], an open-source RAG engine, leverages DeepSeek-R1 to enhance document understanding and enterprise RAG workflows.
@@ II-D2 Mobile Communications
LM agents offer significant advantages for mobile communications by enabling low-cost and context-aware decision-making [65] and personalized user experiences [66]. For instance, NetLLM [65] fine-tunes the LLM to acquire domain knowledge from multimodal data in networking scenarios (e.g., adaptive bitrate streaming, viewport prediction, and cluster job scheduling) with reduced handcraft costs. Meanwhile, NetGPT [66] uses a cloud-edge cooperative LM framework for personalized outputs and enhanced prompt responses in mobile communications via de-duplication and prompt enhancement technologies. ChatNet [75] uses four GPT-4 models to serve as analyzer (to plan network capacity and designate tools), planner (to decouple network tasks), calculator (to compute and optimize the cost), and executor (to produce customized network capacity solutions) via prompt engineering.
@@ II-D3 Intelligent Robots
LM agents play a crucial role in advancing intelligent industrial and service robots [11]. These robots can perform complex tasks such as product assembly, environmental cleaning, and customer service, by perceiving surroundings and learning necessary skills through LMs. In August 2024, FigureAI released Figure 02 [67], a human-like robot powered by OpenAI LM, capable of fast common-sense visual reasoning and speech-to-speech conversation with humans to handle dangerous jobs in various environments. Besides, Unitree unveiled its humanoid robot agent named Unitree G1 [68] in August 2024, which demonstrates the enhanced capabilities brought by LM agents.
@@ II-D4 Autonomous Driving
LM agents are transforming autonomous driving by enhancing vehicle intelligence, improving safety, and optimizing driving experience (e.g., offering personalized in-car experience) [15]. In May 2024, Baidu launched Carrot Run’s sixth-generation unmanned vehicle, built upon the Apollo autonomous driving foundation model (ADFM) [69], an LM agent supporting L4 autonomous driving. Companies such as Tesla, Waymo, and Cruise are also integrating LM agents to enhance autonomous driving systems, aiming for safer and more efficient transportation.